• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name J. Singh

    Log entry time 07:34:00 on May20,2003

    Entry number 103031

    keyword=Target @ Ref Cell, rotated field to 0 deg

    0638 TRANS lasers off, before i turned them off, i noticed the spectrum had some peaks on the low wavelength end. our tune is bad or one or more of the lasers are dying.
    0640 rotating field to 0 deg
    0641 moving to Ref Cell = EMPTY (-016 psig & 000 torr)
    0645 @ ref cell
    0646 @ 0 deg, LONG on, spectrum looks ok